OpenStudy (lilsis76):

Simplifying expressions. How do I do these?? do I convert the -'s to a division looking expression? 1) log(base 10) (x^2 -16) -3 log(base 10) (x+4) +2 log (base 10) x 2) log (base 10) (x^2 -16) - 3 [log (base 10)(x+4) + 2 log (base 10) x] 3) ln(x^3 -1) - ln (x^2 +x +1)

OpenStudy (lilsis76):

@aaronq , do you think you can help explain these to me please. ? I know the log and exponential properties. \[1) \log_{10} (x^{2}-16) -3 \log_{10} (x+4) +2\log_{10} x\] \[2) \log_{10} (x^2-16)-3[\log_{10}(x+4)+2\log_{10}x]\] \[3) \ln (x^3-1)-\ln (x^2+x+1)\]

OpenStudy (lilsis76):

@DebbieG aaronq logged off, do you think you can help me with these problems. ?

OpenStudy (mertsj):

\[\log_{10}m + \log_{10}n =\log_{10}mn \]

OpenStudy (mertsj):

Apply that property to the first and third terms of the first expression.

OpenStudy (lilsis76):

okay,

OpenStudy (mertsj):

You will get: \[\log_{10}x^2(x^2-16) \]

OpenStudy (mertsj):

Now use this property: \[\log_{10}m-\log_{10}n=\log_{10}\frac{m}{n} \]

OpenStudy (lilsis76):

hahaha I got something totally differeent \[\log_{10}x^2 + \log_{10}16 \]

OpenStudy (mertsj):

Why did you change x^2-16 to 16?

OpenStudy (lilsis76):

how did you get an x^2 outside (x^2 -16) ?

OpenStudy (mertsj):